for the balanced chemical reaction C+O2=CO2, we can react 1.0 g of Carbon with exactly 1.0 g of oxygen. Is this statement true or false?
1
Expert's answer
2018-11-28T14:01:09-0500

C + O2 = CO2

According to the balanced chemical reaction: 1mole C – 1 mole O2

Ar (C) = 12, Mr (O2) = 32

n (C) = 1g/12 = 0.0833 moles

n (O2) = 1g/32 = 0.03125 moles

For 1 g of C we need 0.0833 moles*32 = 2.6656 g O2

This statement is false.